OpenStudy (anonymous):

Like four more problems left...I'm getting sleepy. Make "a" the subject of P = 2(a+b)

OpenStudy (anonymous):

P = 2a + 2b (distribution) P - 2b = 2a (P-2b)/2 = a

OpenStudy (anonymous):

lets knock em out

OpenStudy (anonymous):

I'll be posting them on a fresh new thread Sat.

OpenStudy (anonymous):

\[P=2(a+b)\] \[\frac{P}{2}=a+b\] \[\frac{P}{2}-b=a\]

OpenStudy (anonymous):

How are you doing those fractions?

OpenStudy (anonymous):

Jabber lol just click the "equation" button and u can type wat u need to make it clearer

OpenStudy (anonymous):

Here's what it looks like when I do it: \[P/2-b=a\]

OpenStudy (anonymous):

Ehh close enough lol

OpenStudy (anonymous):

I'm wondering how he gets P over 2

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Same here!